At a checkout counter customers arrive at an average of 1.5 per minute. Find the
probabilities that (a) at most 4 will arrive at any given time (b) at least 3 will arrive during
an interval of 2 min time (c) at most 15 will arrive during an interval of 6 min.
1
Expert's answer
2021-04-11T11:48:58-0400

Assume that customers arrival has Poisson distribution with "\lambda = 1.5"

Let "P_k=" the number of arrivals during the given interval.

a)

Let "\lambda_t" be arrival rate during "t" minutes, "\lambda_t=1.5\cdot t";

"P(k\le4)="

"P_0+P_1+P_2+P_3+P_4="

"\sum_0^4{\frac{(1.5t)^k}{k!}e^{-1.5t}}"


b)

"\lambda=1.5\cdot 2=3"

"P(k\ge3)=1-P_0-P_1-P_2="

"1-\frac{3^0}{0!}\cdot e^{-3}-\frac{3^1}{1!}\cdot e^{-3}-\frac{3^2}{2!}\cdot e^{-3}="

"1-e^{-3}(1+3+9/2)=1-\frac{17}{2}e^{-3}\approx 0.577."

c)


"\lambda=1.5\cdot6=9"

Use Normal distribution "N(\lambda,\lambda)" to approximate Poisson distribution:

"F_{poisson}(x,\lambda)\approx F_{normal}(x+1/2,\lambda,\lambda)"

"F_{poisson}(x,9)\approx F_{normal}(x+1/2,9,9)"

"F_{normal}(x+1/2,9,9)=Pr(X_{normal}\le15.5)="

"Pr(\frac{X_{normal}-9}{3}\le\frac{15.5-9}{3})=Pr(Z\le2.1667)\approx0.985"

Another option is to use cumulative Poisson distribution table, this method gives 0.978.

Answer: a) "\sum_0^4{\frac{(1.5t)^k}{k!}e^{-1.5t}}", b) 0.423, c) 0.978.